Please find and solve x in the following equation, showing steps:
2x-1/3 - 6x-2/5 = 1/3

Answer by Cromlix(4381)

2x-1/3 - 6x-2/5 = 1/3
Common denominator = 15
10x -5 - (18x - 6) = 5
10x - 5 - 18x + 6 = 5
-8x = 5 - 1
-8x = 4
x = 4/-8
x = -1/2
Proof:
2x-1/3 - 6x-2/5 = 1/3
(2(-1/2) -1)/3 - (6(-1/2) -2)/5 = 1/3
-2/3 - (-5/5) = 1/3
-2/3 + 5/5 = 1/3
Common denominator = 15
-10 + 15 = 5
Hope this helps.
:-)